I am having a few issues if someone could help. The first problem is
3 with exponent of x+4 = 7 with exponent of 2x-1
Through all my notes I am finding nothing like this, my exmples are where I would take 3 as my base for 2nd equation

Answer by stanbon(75887)

3 with exponent of x+4 = 7 with exponent of 2x-1
----------------
3^(x+4) = 7^(2x-1)
---------
Since the variable is in the exponent, take the log::
(x+4)*log(3) = (2x-1)*log(7)
-------------------------------------
Now, get the x terms together::
x(log3-2*log(7)) = log(7)-4log(3)
x(-1.213) = -1.063
x = 0.877
-------------
Cheers,
Stan H.
--------------